# The Loss of Control Playbook: Degrees, Dynamics, and Preparedness

### v3

@@
Figure 1: Distribution of 12 concrete LoC scenarios.
@@
  • Deviation: captures events that cause some harm or inconvenience but lack the requisite severity and persistence to reach the economic consequences threshold that the U.S. Department of Homeland Security, federal agencies, and the intelligence community use to demarcate national-level events in the Strategic National Risk Assessment.
@@ Introduction
Notwithstanding the rising interest, the topic area of LoC itself is nascent and lacks a common, clear definition. In this context, we found that neither AI literature (see Section 1.A.1.) nor other safety-critical industries (see Section 1.A.1.) provide sufficient support for a common, clear, and, in turn, widely actionable definition of LoC. The absence of such a definition may soon contribute to epistemic confusion around LoC, especially as more decision- and policymakers engage with the topic. In turn, this could make the identification of early warning shots challenging and undermine the shared effort by the aforementioned and future legal frameworks to target and mitigate LoC threats appropriately. Finally, an absence of clarity could also lead to ‘crying wolf’ situations for phenomena that fall short of LoC, and ultimately contribute to fostering skepticism of LoC itself as a remote, impalpable, and speculative threat.
@@
  • A straightforward framework to minimize the threat of LoC today, which works around existing uncertainties regarding LoC-relevant capabilities and propensities, and that decision- and policymakers can, therefore, leverage immediately (see Chapter 2).
@@
In turn, this allowed us to extrapolate some defining elements of LoC scenarios and to support decision- and policymakers in carving out the content and boundaries of LoC more clearly. Our analysis further enabled us to assess why LoC definitions across the literature may differ and to propose a novel conceptualization of LoC. Specifically, in our literature review, we found that LoC scenarios often differ in terms of their severity and persistence. These two metrics broadly capture intuitions around, among other things, the number of people affected, the degree to which those people are affected, and the difficulty of stopping the LoC incident. We subsequently used these metrics to plot concrete LoC scenarios from the literature (see footnote 15) onto an experimental graph, allowing us to provide additional context to existing descriptions of LoC.
@@ 1.A.2 Findings from Loss of Control Scenarios in Literature
First, we sought to contextualize the area indicated in the lower left of our graph (see footnote 15). In doing so, we came across the economic consequence thresholds that the U.S. Department of Homeland Security (DHS), alongside federal agencies and the intelligence community, identified as “necessary to create a national level-event” in the Strategic National Risk Assessment (SNRA) [DHS_SNRA_2011]. Specifically, in 2011, DHS led an effort to “identify the types of incidents that pose the greatest threat to the Nation’s homeland security,” including various natural, technological, and adversarial, human-caused hazards [DHS_SNRA_2011]. [^16] We decided to adopt the SNRA as a term of reference because the U.S. is the default location where the majority of literature assumes LoC scenarios to occur. For instance, ‘Scenario 1’ in [Kalra_GeopoliticsAGI] is an example of a scenario that envisages AI-caused electricity blackouts starting in the U.S. In some cases, the location of the LoC scenario described by literature is ambiguous (for instance, ‘Scenario 4’ in [Kalra_GeopoliticsAGI]), so for the purposes of making a comparable economic impact estimate, we assumed that the scenario occurred within the U.S. to ensure consistency across scenarios. For example, earthquakes, floods, hurricanes, wildfires, and cyberattacks against physical infrastructure meet the threshold if they result in direct economic losses exceeding \$100 million [DHS_SNRA_2011]. Similarly, a cyberattack against data [^17] In this context, data refers to the information contained in a computer system or data processes. is considered a national-level event if it results in economic losses of \$1 billion or more. In this research report, we adopt this last threshold set by the SNRA at \$1 billion or greater (adjusting it for inflation to approximately \$1.4 billion) to provide some contextualization of LoC scenarios (see footnote 15). We refer to this threshold, i.e., the \$1.4 billion, for convenience, as the ‘national risk assessment threshold.’ In summary, the national risk assessment threshold is the highest threshold based on economic consequences that DHS, federal agencies, and the intelligence community identified in the Strategic National Risk Assessment to demarcate a national-level event (i.e., a threat or hazard that has the potential to significantly impact the U.S. homeland security). We indicate this threshold in orange in Figure 1.
@@ 2.A Capabilities and Propensities: Intuitive but not Actionable Today
2.B. Focus on Deployment Environment, Affordances, and Permissions
@@ 2.B.1 Deployment Context
  • Second, in both cases (normal deployment context and high-risk deployment context) to assess the potential for cascading failures across interconnected systems, including both AI- and non-AI systems, through, for instance, threat modeling and red teaming [Anthropic2025resp, shevlane2023modelevaluationextremerisks, koessler2024riskthresholdsfrontierai], and to consider whether overall risk can be limited by applying 2.B.2. and 2.B.3.282828We note that deployment in select high-stakes deployment contexts (for instance, direct control of nuclear weapons), where LoC would lead to critical and absolute consequences, should be rejected regardless of the potential benefits and safeguards.
@@ 2.C Reflections
In the longer term, the DAP framework is likely to face two main challenges. First, it may not always be desirable for society to limit deployment contexts, affordances, and permissions in the way we described, since these interventions simultaneously limit what the AI system can actually do, therefore reducing its potential usefulness. Indeed, as AI systems become more capable across complex use cases, there may be growing economic and strategic incentives to deploy them with a wide range of affordances and permissions (including in high-risk environments) with promises of significant benefits, such as efficiency increases and obtaining or maintaining a strategic advantage (including on the geopolitical level). Second, it is plausible that future significantly more advanced AI systems will be able to meaningfully persuade or compel humans to concede more affordances and permissions [RAND_LOC_Response_Planning, RAND_Preparedness, dassanayake2025manipulationattacksmisalignedai] or even independently circumvent the blockers established through the DAP framework in a way we cannot currently foresee.
@@
3.A. The Pathway to a State of Vulnerability
@@ 3.B. The Catalyst Triggering LoC
3.C. Implications of Living with a State of Vulnerability
@@
If (2.i) occurs, then a LoC outcome has materialized. Since we are referring to LoC to capture Bounded LoC and Strict LoC (see p.19), this materialization implies that society finds itself faced with an outcome in either of these two categories. Depending on the severity of the Bounded LoC, which will be hard to predict ex ante, society may be willing to incur substantial costs to avoid such an outcome. Certainly, society will wish to avoid the materialization of Strict LoC, which poses an irreversible situation with no scope left for action. In both instances, branch (2.i) implies that a LoC outcome has occurred, and therefore, this is the outcome to avoid at high or all costs.

### v4

@@
Figure 1: This graph plots 12 concrete LoC scenarios identified in the literature. We utilized economic impact as a proxy for severity and persistence, which are both mapped on arbitrary axes of 0-100. These data points inform our proposed three-part taxonomy, splitting apart Deviation, Bounded LoC, and Strict LoC. By sharpening the conceptual boundaries of LoC, this taxonomy helps decision-makers to understand the different degrees of LoC and hence better prioritise between risk-reduction strategies. See Section 1.A.2 for more detail on the methodology behind and limitations of this visualization.
@@
  • Deviation: captures events that cause some harm or inconvenience but lack the requisite severity and persistence to reach the economic consequences threshold that the the U.S.’s Department of Homeland Security, Intelligence Community, and other components use to demarcate national-level events in the Strategic National Risk Assessment.
@@ Introduction
Notwithstanding the rising interest, the topic area of LoC itself is nascent and lacks a common, clear definition. In this context, we found that neither AI literature nor other safety-critical industries (see Section 1.A.1) provide sufficient support for a common, clear, and, in turn, widely actionable definition of LoC. The absence of such a definition may soon contribute to epistemic confusion around LoC, especially as more decision- and policymakers engage with the topic. In turn, this could make the identification of early warning shots challenging and undermine the shared effort by the aforementioned and future legal frameworks to target and mitigate LoC threats appropriately. Finally, an absence of clarity could also lead to ‘crying wolf’ situations for phenomena that fall short of LoC, and ultimately contribute to fostering skepticism of LoC itself as a remote, impalpable, and speculative threat.
@@
  • A straightforward framework to minimize the threat of LoC today, which works around existing uncertainties regarding LoC-relevant capabilities and propensities, and that decision- and policymakers can, therefore, leverage immediately (see .
@@ 1.A.2 Findings from Loss of Control Scenarios in Literature
First, we sought to contextualize the area indicated in the lower left of our graph (see footnote 15). In doing so, we came across the economic consequence thresholds that the U.S.’s Department of Homeland Security (DHS), the Intelligence Community, and other components, identified as “necessary to create a national level-event” in the Strategic National Risk Assessment (SNRA) [DHS_SNRA_2011]. Specifically, in 2011, DHS led an effort to “identify the types of incidents that pose the greatest threat to the Nation’s homeland security,” including various natural, technological, and adversarial, human-caused hazards [DHS_SNRA_2011]. [^16] We decided to adopt the SNRA as a term of reference because the U.S. is the default location where the majority of literature assumes LoC scenarios to occur. For instance, ‘Scenario 1’ in [Kalra_GeopoliticsAGI] is an example of a scenario that envisages AI-caused electricity blackouts starting in the U.S. In some cases, the location of the LoC scenario described by literature is ambiguous (for instance, ‘Scenario 4’ in [Kalra_GeopoliticsAGI]), so for the purposes of making a comparable economic impact estimate, we assumed that the scenario occurred within the U.S. to ensure consistency across scenarios. For example, earthquakes, floods, hurricanes, wildfires, and cyberattacks against physical infrastructure meet the threshold if they result in direct economic losses exceeding \$100 million [DHS_SNRA_2011]. Similarly, a cyberattack against data [^17] In this context, data refers to the information contained in a computer system or data processes. is considered a national-level event if it results in economic losses of \$1 billion or more. In this research report, we adopt this last threshold set by the SNRA at \$1 billion or greater (adjusting it for inflation to approximately \$1.4 billion) to provide some contextualization of LoC scenarios (see footnote 15). We refer to this threshold, i.e., the \$1.4 billion, for convenience, as the ‘national risk assessment threshold.’ In summary, the national risk assessment threshold is the highest threshold based on economic consequences that DHS, the Intelligence Community and other components identified in the Strategic National Risk Assessment to demarcate a national-level event (i.e., a threat or hazard that has the potential to significantly impact the U.S. homeland security). We indicate this threshold in blue in Figure 1.
@@ 2.A Capabilities and Propensities: Intuitive but not Actionable Today
2.B Focus on Deployment Environment, Affordances, and Permissions
@@ 2.B.1 Deployment Context
  • Second, in both cases (normal deployment context and high-stakes deployment context) to assess the potential for cascading failures across interconnected systems, including both AI- and non-AI systems, through, for instance, threat modeling and red teaming [Anthropic2025resp, shevlane2023modelevaluationextremerisks, koessler2024riskthresholdsfrontierai], and to consider whether overall risk can be limited by applying 2.B.2 and 2.B.3.282828We note that deployment in select high-stakes deployment contexts (for instance, direct control of nuclear weapons), where LoC would lead to critical and absolute consequences, should be rejected regardless of the potential benefits and safeguards.
@@ 2.C Reflections
In the longer term, the DAP framework is likely to face two main challenges. First, it may not always be desirable for society to limit deployment contexts, affordances, and permissions in the way we described, since these interventions simultaneously limit what the AI system can actually do, therefore reducing its potential usefulness. Indeed, as AI systems become more capable across complex use cases, there may be growing economic and strategic incentives to deploy them with a wide range of affordances and permissions (including in high-stakes environments) with promises of significant benefits, such as efficiency increases and obtaining or maintaining a strategic advantage (including on the geopolitical level). Second, it is plausible that future significantly more advanced AI systems will be able to meaningfully persuade or compel humans to concede more affordances and permissions [RAND_LOC_Response_Planning, RAND_Preparedness, dassanayake2025manipulationattacksmisalignedai] or even independently circumvent the blockers established through the DAP framework in a way we cannot currently foresee.
@@
3.A The Pathway to a State of Vulnerability
@@ 3.B The Catalyst Triggering LoC
3.C Implications of Living with a State of Vulnerability
@@
If (2.i) occurs, then a LoC outcome has materialized. Since we are referring to LoC to capture Bounded LoC and Strict LoC (see Section 1.A.2.b), this materialization implies that society finds itself faced with an outcome in either of these two categories. Depending on the severity of the Bounded LoC, which will be hard to predict ex ante, society may be willing to incur substantial costs to avoid such an outcome. Certainly, society will wish to avoid the materialization of Strict LoC, which poses an irreversible situation with no scope left for action. In both instances, branch (2.i) implies that a LoC outcome has occurred, and therefore, this is the outcome to avoid at high or all costs.